At 23m below the surface of the sea (density of water= 1,014 kg/m3), where the temperature is 4.6 Celsius, a diver exhales an air bubble having a volume of 2.0 cm3. If the surface temperature of the sea is 21 Celsius and the atmospheric pressure measured at the surface is 101,500 Pa, what is the volume of the bubble just before it breaks the surface? Take g=9.8 m/s2 and the absolute zero to be at -273 Celsius. State your answer in cm3 to the nearest 0.01 cm3.


1
Expert's answer
2021-10-03T13:24:58-0400
"V_f=V_i\frac{T_f}{T_i}\frac{P+\rho gh}{P}\\
V_f=2\frac{294}{277.6}\frac{101500+(1014)(9.8)(23)}{101500}=6.89\ cm^3"